OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 2 and contains the point (3, 7).

OpenStudy (bahrom7893):

y - 7 = 2(x-3) y - 7 = 2x - 6 y = 2x-6+7 y = 2x+1